OpenStudy (anonymous):

prove the integer 53^103+103^53 is divisible by 39

OpenStudy (anonymous):

Are you familiar with modular arithmetic? (I hope).

OpenStudy (anonymous):

no... :( ...please show me the steps

OpenStudy (anonymous):

Oh.

myininaya (myininaya):

how do you do this without number theory?

OpenStudy (anonymous):

That's what I'm trying to think about.

myininaya (myininaya):

what class is this?

myininaya (myininaya):

If you can tell me the class, maybe I can come up with a way to help you that you will understand

OpenStudy (anonymous):

I'm in 10th...just saw this problem and eager to know how to do it

myininaya (myininaya):

bond what was your modular arithmetic way?

OpenStudy (anonymous):

Well \(53\equiv14\) while \(103\equiv-14\), but that's as far as I got.

OpenStudy (anonymous):

Plus 53 is an odd power, so the - stays.

myininaya (myininaya):

oh wait 103 is congruent to 25 and 25+14=39

myininaya (myininaya):

whenever i took number theory we wrote the numbers could only be between 0 and whatever number minus 1 we are writing are congruences in

OpenStudy (anonymous):

Yeah, it's the same idea.

OpenStudy (anonymous):

\(-14\equiv25\)

myininaya (myininaya):

your right though

OpenStudy (anonymous):

Ok, here's the deal: \[53^{103}\equiv14^{103}=(14^2)^{51}*14\equiv1^{51}*14\equiv14({\rm mod} 39)\]

myininaya (myininaya):

14 is congruent to 1 mod 39?

OpenStudy (anonymous):

no, but 14^2 is

OpenStudy (anonymous):

14^2=196 and 195 is divisible by 39

myininaya (myininaya):

jaebond I think I might have to become your fan. I like your number theory skills.

OpenStudy (anonymous):

Why thank you, I've been brushing up lately.

myininaya (myininaya):

Its been awhile since i had number theory

OpenStudy (anonymous):

So maria, the idea behind modular arithmetic is that we say that two numbers are the same if they differ by a multiple of modulus. In this case, if the "modulus" is 39, we say that 14 and 53 are the same. We also say that 92 is the same and we can write \(14\equiv53\equiv92 ({\rm mod }39)\)

OpenStudy (anonymous):

Most of the properties of the integers hold as normal, except that you usually can't divide.

OpenStudy (anonymous):

So the first step from what I did earlier looks like \[53^{103}=53*53*53*...*53\] where there are 103 53's. But \(53\equiv14\), so this becomes \(14*14*14*...*14=14^{103}\).

OpenStudy (anonymous):

Does this make sense maria?

OpenStudy (anonymous):

hmmmm.....that's clear enough.... thanks both of you.... :)

OpenStudy (anonymous):

why is 103 congruent to 25 ?

OpenStudy (anonymous):

Do you know why this is working, Maria?

OpenStudy (anonymous):

I would start off by realising 39 is divisible by 3 and 13, then show that 53^103 + 1-3^53 is divisible by both 3 and 13. It will then be divisible by 39.

OpenStudy (anonymous):

So, 53mod3 == -1 (which means you can write 53 as 3n-1, for some integer n). So 53^103 = 3k-1, for some integer k. Similarly, 103mod3==1, which means 103^53 can be written 3p+1, for some integer p. When you add them, 3k-1+3p+1 = 3(k+p)=3j, for some integer j. Therefore, 3 divides (103^53 + 53^103). That's the first part.

OpenStudy (anonymous):

103mod13==-1, so 103^53 = 13s-1, for some integer s. 53mod13==1, so 53^13= 13t+1, for some integer t. Adding them, 103^53+53^13=13(s+t)=13r, for some integer r. Therefore, 13 divides your sum. Since your sum is divisible by 3 and 13, and 3 and 13 are prime, it's divisible by 39.

OpenStudy (anonymous):

ohhhhhhhh.....I see